OpenStudy (anonymous):

How do you solve X+6=3(5-x)?

OpenStudy (anonymous):

x+6 = 3(5-x) x+6 = 15-3x x+3x = 15-6 4x = 9 x = 9/4

OpenStudy (anonymous):

X+6=15-3x x+3x=15-6 4x=9 x=9/4= 2 1/4

OpenStudy (anonymous):

Oh okay I get it Thank you!